# Recent Results on the Tetraneutron

### v1

@@ 2.5 Photodisintegration of 209Bi in Yerevan, Armenia
A team at the Alikanyan National Science Laboratory in Yerevan, Armenia [9] used Bremsstrahlung from electron beams to disintegrate 209Bi. After irradiation of a pure 209Bi target with photons they measured the -delayed -radiation and the corresponding decay curves of the two most abundant transitions populated in the /EC-decay of 205Bi (T d [10]). At a nominal electron energy of 30 MeV - that is 0.50 MeV above the threshold for the 209Bi(Bi reaction - they observed a cross section, weighted with the Bremsstrahlung spectrum, of mb[11]. At 28 MeV, however, the cross section is negligible: . They compared the 30 MeV cross section for the emission of four neutrons with calculations using the TALYS1.9 and FLUKA models taking into account the energy spread of the electron beam of , with the result that the calculations underestimate the observed cross section at 30 MeV beam energy by more than an order of magnitude. They concluded that a correlated emission of four neutrons might explain this discrepancy because it is not included in the models. In their publication [9] they propose, that the observed resonance structure from [8] could be a good explanation, if this is not caused by the reaction mechanism as postulated by Lazauskas et al. [12]. However, if we plot into their calculated photon spectrum the resonance curve measured by Duer et al. [8] (see Fig. 1) we note only a minuscule overlap, even with the photon spectrum from 30 MeV electrons. A much better overlap of the spectrum from 30 MeV electrons can be observed with the bound tetraneutron claimed by Faestermann et al. [6].
@@ 4 Synthesis of all experimental results
Here we want to try to understand the presently available experimental information. In Fig. 2 we have summarized the results of the experiments described above with respect to the observed width and energy available for the decay into four neutrons. Only the experiment of Kotanjyan et al. [9] cannot specify the energy, but as it may be concluded from Fig. 1 a state very close below or above threshold seems to explain best the large cross section difference between electron energies of 28 and 30 MeV. The result of Marqués et al. [2, 3] is only an upper limit for the energy and, as the results of Muzalevskii et al. [13], gives no information on the width. Kisamori et al. [5] and Faestermann et al. [6] have upper limits for the width, whereas Duer et al. [8] has also a lower limit for the width. Concerning the energies Duer et al. [8] and Muzalevskii et al. [13] are compatible with a weighted mean of MeV. All three results are certainly not compatible with the very precise, negative, value of Faestermann et al. [6]. The values of Marqués et al. [2, 3] and Kisamori et al. [5] may be compatible with both. Thus, Faestermann et al., on the one hand, and Duer et al. and Muzalevskii et al., on the other, cannot have observed the same state of the tetraneutron. However, could it not be that Faestermann et al. have observed the ground state and Duer et al. and Muzalevskii et al. the first excited state? We would expect the ground state to be formed from two neutrons in the shell and two neutrons in the shell coupled to spin zero. The first excited state would then have the two neutrons coupled to . The energy difference would be MeV. For the four even-even neighbours with : 6He, 8Be, 10C and 12O the average excitation energy of the -state is MeV(RMS). (The uncertainty is the root-mean-square deviation.) In the target nucleus 7Li of the experiment of Faestermann et al.[6] the neutrons are already in the lowest configuration, its spin is only determined by the protons which are removed in the reaction. In addition, low energy transfer reactions do not rely on a sudden process where the structural overlap of initial and final states are especially important but go through a relatively long compound phase mixing single particle wave functions. However, in the -knockout as well as in the -transfer reactions it can be assumed that the is formed from two neutrons in the shell which are removed. Therefore, only neutrons would remain and it seems quite likely that the four neutrons in the final channel are in a state. We have argued in a similar way before [31]. Or, to argue with radial dimensions, the 8He matter radius is with 2.49(4) fm [32] considerably larger than the charge radius of 1.959(16) fm [33, 34] (Note that the charge radius of 4He is only 1.681(4) fm [35], but in 8He it can move around the common center of gravity of 4He and the four extra neutrons.). That means that the additional four neutrons form a halo around the strongly bound -particle. When the latter is removed there may be only a small overlap between the halo and the ground state configuration of the tetraneutron. Duer at al.[8] implemented the 8He wave function in the initial state and derived a quite large overlap with the configuration needed for the knockout. Nevertheless, this does not allow a conclusion on the preferred configuration that will be populated.
@@ 5 Conclusions
Acknowledgements We are grateful to Horst Lenske for a critical reading of the manuscript.

### v2

@@ 2.5 Photodisintegration of 209Bi in Yerevan, Armenia
A team at the Alikanyan National Science Laboratory in Yerevan, Armenia [9] used Bremsstrahlung from electron beams to disintegrate 209Bi. After irradiation of a pure 209Bi target with photons they measured the -delayed -radiation and the corresponding decay curves of the two most abundant transitions populated in the /EC-decay of 205Bi (T d [10]). At a nominal electron energy of 30 MeV - that is 0.50 MeV above the threshold for the 209Bi(Bi reaction - they observed a cross section, weighted with the Bremsstrahlung spectrum, of . At 28 MeV, no signal of the 205Bi production was obseved, with an upper limit of the weighted cross section of . They compared the 30 MeV cross section for the emission of four neutrons with calculations using the TALYS1.9 and FLUKA models taking into account the energy spread of the electron beam of , with the result that the calculations underestimate the observed cross section at 30 MeV beam energy by more than an order of magnitude. They concluded that a correlated emission of four neutrons might explain this discrepancy because it is not included in the models. In their publication [9] they propose, that the observed resonance structure from [8] could be a good explanation, if this is not caused by the reaction mechanism as postulated by Lazauskas et al. [11]. However, if we plot into their calculated photon spectrum the resonance curve measured by Duer et al. [8] (see Fig. 1) we note only a minuscule overlap, even with the photon spectrum from 30 MeV electrons. A much better overlap of the spectrum from 30 MeV electrons can be observed with the bound tetraneutron claimed by Faestermann et al. [6].
@@ 4 Synthesis of all experimental results
Here we want to try to understand the presently available experimental information. In Fig. 2 we have summarized the results of the experiments described above with respect to the observed width and energy available for the decay into four neutrons. Only the experiment of Kotanjyan et al. [9] cannot specify the energy, but as it may be concluded from Fig. 1 a state very close below or above threshold seems to explain best the cross section difference between electron energies of 28 and 30 MeV. The result of Marqués et al. [2, 3] is only an upper limit for the energy and, as the results of Muzalevskii et al. [12], gives no information on the width. Kisamori et al. [5] and Faestermann et al. [6] have upper limits for the width, whereas Duer et al. [8] has also a lower limit for the width. Concerning the energies Duer et al. [8] and Muzalevskii et al. [12] are compatible with a weighted mean of MeV. All three results are certainly not compatible with the very precise, negative, value of Faestermann et al. [6]. The values of Marqués et al. [2, 3] and Kisamori et al. [5] may be compatible with both. Thus, Faestermann et al., on the one hand, and Duer et al. and Muzalevskii et al., on the other, cannot have observed the same state of the tetraneutron. However, could it not be that Faestermann et al. have observed the ground state and Duer et al. and Muzalevskii et al. the first excited state? We would expect the ground state to be formed from two neutrons in the shell and two neutrons in the shell coupled to spin zero. The first excited state would then have the two neutrons coupled to . The energy difference would be MeV. For the four even-even neighbours with : 6He, 8Be, 10C and 12O the average excitation energy of the -state is MeV(RMS). (The uncertainty is the root-mean-square deviation.) In the target nucleus 7Li of the experiment of Faestermann et al.[6] the neutrons are already in the lowest configuration, its spin is only determined by the protons which are removed in the reaction. In addition, low energy transfer reactions do not rely on a sudden process where the structural overlap of initial and final states are especially important but go through a relatively long compound phase mixing single particle wave functions. However, in the -knockout as well as in the -transfer reactions it can be assumed that the is formed from two neutrons in the shell which are removed. Therefore, only neutrons would remain and it seems quite likely that the four neutrons in the final channel are in a state. We have argued in a similar way before [30]. Or, to argue with radial dimensions, the 8He matter radius is with 2.49(4) fm [31] considerably larger than the charge radius of 1.959(16) fm [32, 33] (Note that the charge radius of 4He is only 1.681(4) fm [34], but in 8He it can move around the common center of gravity of 4He and the four extra neutrons.). That means that the additional four neutrons form a halo around the strongly bound -particle. When the latter is removed there may be only a small overlap between the halo and the ground state configuration of the tetraneutron. Duer at al.[8] implemented the 8He wave function in the initial state and derived a quite large overlap with the configuration needed for the knockout. Nevertheless, this does not allow a conclusion on the preferred configuration that will be populated.
@@ 5 Conclusions
Acknowledgements We are grateful to Horst Lenske and Tigran Kotanjyan for a critical reading of the manuscript.